(13%) Problem 5: ms. A rock takes 2.65 s to ht the ground when it is thrown straight up from the cliff with an initial velocity of 7 IS 15 50% Part (a) Calculate the height of the cliff in m. pleted-14.005 pleted h=14.01 V Correct! leted pleted | tial | 50% Part (b) How long would it take to reach the ground if it is thrown straight down with the same speed? Grade Summary Deductions Potential 91 cotan() | asm() 1 acos() | atanOacotansinh0 coshO tanhO cotanho Submissions Atempts remaining 0% per attempt) detailled view 0 END 3% Degrees Radians NO BACKSPACE CLEAR SubmitHtFeedback Igive up! 0% deduction per Hints:096s deduction per hint. Hints remaining: 1 All cotent 2019 Expert TA LLC


I need to know all the steps to get parts a+b. thank you.
Answer #1

1.

S = ut + 1/2 g t^2
Here
u = 7.7 m/s
t= 2.65 sec
So
S= 7.7*2.65 + 1/2*(-9.8)*2.65^2
S = - 14.005 meter
Or
14.005 meter downward <------- answer

2.
From previous equation
14.005 = 7.7t + 1/2(9.8)*t^2

4.9t^2 + 7.7t - 14.005 = 0
On solving
t = -2.65 sec NEGLECTING NEGETIVE

t = 1.07 sec <--------- ANSWER